When parking a car in a downtown parking lot, drivers pay according to the number of hours or fraction thereof. The probability distribution of the number of hours cars are parked has been estimated as follows:

X 1 2 3 4 5 6 7 8
P(X) 0.224 0.142 0.106 0.08 0.057 0.039 0.033 0.319

A. Mean =

B. Standard Deviation =

The cost of parking is 2.25 dollars per hour. Calculate the mean and standard deviation of the amount of revenue each car generates.

A. Mean =

B. Standard Deviation =
Answer #1

Answer

(A)Mean =

using the data table

And

(B)Standard deviation =

x(bar) is mean = 4.448, P(X) is probability and X is the given values in the data table

this implies

The cost of parking is 2.25 dollars per hour

(A) mean =

= 4.448*2.25

= $10.008

(B) Standard deviation =

= 8.10*2.25

= $18.225